# Advanced safety filter based on SOS Control Barrier and Lyapunov Functions

### v2

@@ Abstract
This paper presents a novel safety filter framework that ensures both safety and the preservation of the legacy control action within a nominal region. This modular design allows the safety filter to be integrated into the control hierarchy without compromising the performance of the existing legacy controller within the nominal region. This is accomplished by formulating multiple Control Barrier Functions (CBFs) and Control Lyapunov-like Functions (CLFs) conditions, alongside a forward invariance condition for the legacy controller, as sum-of-squares constraints utilizing Putinar’s Positivstellensatz. Additionally, the state-dependent inequality constraints of the resulting Quadratic Program – encoding the CBF and CLF conditions – are designed to remain inactive within the nominal region, ensuring perfect tracking of the legacy control action. Our safety filter design is also the first to include quadratic input constraints, and does not need an explicit specification of the attractor, as it is implicitly defined by the legacy controller. To avoid the chattering effect and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the Quadratic Program are selected to be sufficiently regular. Finally, we demonstrate the method in a detailed case study involving the control of a three-phase ac/dc power converter.
@@ I Introduction
In various engineering applications, it is of much interest to endow a well-performant controller with safety guarantees by minimally changing its control action. This concept is commonly referred to as a safety filter, and a comprehensive overview of the subject can be found in [1]. An illustrative example is a finely-tuned control design based on linearization, which must be extended to adhere to hard state constraints in the nonlinear closed-loop system. We refer to such a well-performant controller as the legacy controller. A popular type of safety filters relies on solving a Quadratic Program (QP) at every state using a Control Barrier Function (CBF). By continuously solving the QP, the control action of the legacy controller is smoothly adjusted as the system approaches the boundary of the safe set $\mathcal{X}_{s}$. The main challenges in deploying such a safety filter lies in finding a viable CBF candidate and adhering to input constraints.
Even though safety filters based on a QP are well-established, they are frequently not adopted in current industrial setups due to their risk of altering the legacy control behavior during nominal operations. In practice, the legacy control parameters are often meticulously tuned and thoroughly evaluated, typically involving extensive hardware-in-the-loop simulations, which require significant time and resources. As a result, the safety filter is expected to intervene only when the system approaches state constraints, meaning when it is leaving the nominal operating regime.
@@
This paper presents a safety filter concept that guarantees the preservation of the legacy control action within a nominal region $\mathcal{X}_{n}$, which contains the desired closed-loop equilibrium point (or another attractor), as depicted in Figure 1. Unlike in a reach-avoid problem [2], the nominal region $\mathcal{X}_{n}$ does not represent a predetermined set of target states. Instead, it is selected during the design process (often at an earlier stage) to encompass the nominal operating regime of the legacy controller while maintaining a sufficiently large transitional region for a smooth safety filter operation. The specifications are inherently more stringent than those of a standard safety filter, but may therefore be deployed in an industrial setup without interference in the original line-up.
Our safety filter is realized based on solving a Quadratic Constraint Quadratic Program (QCQP) at every state using multiple CBFs and Control Lyapunov-like Functions (CLFs). A QCQP instead of a simple QP is employed to encode quadratic input constraints, which are common, e.g., in power electronic applications. While CBF conditions (encoded as linear constraints of the QCQP) are used to ensure safety w.r.t. $\mathcal{X}_{s}$, we leverage the CLF condition (also encoded as linear constraints of the QCQP) to specify a dissipation rate ensuring a finite-time convergence from the boundary of the safe set to the nominal region $\mathcal{X}_{n}$. The finite-time convergence property is not intended to enhance control performance but to ensure a rapid return to the nominal operating regime following an unforeseen event. We ensure that the nominal set contains the desired attractor by selecting it forward invariant with respect to the legacy controller. For identifying compatible CBF and CLF candidates that adhere to the linear and quadratic input constraints, as well as a solution for the QCQP at every state, we propose a Sum-of-Squares (SOS) optimization approach. Preserving the legacy control action is achieved by ensuring that the inequality constraints of the QCQP remain inactive within $\mathcal{X}_{n}$. This necessitates taking into account the legacy controller when searching for appropriate CBF and CLF candidates. To solve the resulting SOS program, we deploy an alternating algorithm that simultaneously searches for a feasible controller in the class of rational functions of the state.
@@
The concept of our safety filter resembles the integration of a CLF with a CBF to attain both safety and stabilization. Such a unification, although just with a single function, has been initially proposed in [3]. The integration of both CLF and CBF into an optimization-based controller has since become a well-established concept, see the survey in [4]. Feasibility of such an optimization-based controller can be accomplished by introducing a slack variable that deactivates CLFs and CBFs based on their respective priorities, as shown in [5]. Alternatively, it is possible to compute compatible CLFs and CBFs using SOS tools, ensuring that for every state, there exists an input $u$ satisfying all CBF and CLF conditions. In order to encode compatibility, there is one approach that introduces an SOS constraint for each CLF/CBF pair in the absence of inputs constraints [6]. The second approach establishes compatibility by introducing a polynomial or rational state-feedback controller, as outlined in [7, 8]. Such a controller can also be used to enforce linear input constraints, as demonstrated in [9]. Another noteworthy approach for imposing input constraints using a contrapositive statement, albeit without ensuring compatibility between CBFs and CLFs, is introduced in [10]. These approaches introduce auxiliary polynomials that render the resulting SOS problem bilinear in its decision variables. In [11], an alternating algorithm is proposed to solve this bilinear SOS problem. One challenge in seeking CLF and CBF candidates through an alternating algorithm is to identify an initial feasible point. In their work in [12], the authors constrained their search for an initial CBF to scalar functions of a specific form. This restriction allows to reformulate the problem with linearity in its decision variables.
After our previous work [8] on synthesizing a rational controller to find compatible CLF and multiple CBFs using SOS optimization, and through collaboration with industrial partners, we identified a recurring issue: any modifications to the existing legacy control architecture should not compromise the nominal control performance. A prime example for this is the integration of a safety filter. In this paper, we propose a novel way of implementing a safety filter without disturbing the legacy controller within the nominal region. The key advancements are summarized as follows:
  • First, we propose a novel safety filter implementation that, by design, ensures finite-time convergence to the nominal region $\mathcal{X}_{n}$ and perfect tracking of the legacy control action within that region.
@@
  • Second, multiple CBFs are used to the increase flexibility of the resulting safe set. Instead of imposing each CBF condition separately, we relax these conditions when the corresponding CBF is not active, similar to the idea in [13]. Moreover, the CLF condition can also be relaxed by leveraging the forward invariance property of the CBFs, similar to [14] for autonomous systems.
  • Third, the desired attractor (e.g., an equilibrium point or limit cycle), which is stabilized by the legacy controller in the nominal region, does not need to be specified explicitly, as it is implicitly specified through the legacy controller as a forward invariance condition.
@@
  • Fourth, we introduce a novel approach for encoding linear and quadratic input constraints on the rational controller. This ensures feasibility of the resulting QCQP-based controller, even in case of quadratic input constraints – an issue that, to the authors’ knowledge, is unsolved.
@@
To avoid chattering and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the QCQP are selected to be sufficiently regular. Finally, we tested our framework on a detailed three-phase ac/dc converter system connected to an infinite bus, where we successfully synthesized a safety filter that ensures satisfaction of the state constraints given by the dc-link voltage and line current, while adhering to quadratic input constraints. We further tracked an a priori unspecified quadrature current reference of the ac/dc converter by integrating the reference as a state into a slightly modified system model.
@@ II-A Notation
The shorthand $\mathcal{I}_{t}:=\{1,2,...,t\}$ designates a range of natural numbers from 1 to $t$. $R[x]$ refers to the set of scalar polynomials in variables $x\in\mathbb{R}^{n}$, and $\Sigma[x]$ refers to the set of scalar SOS polynomials in variables $x\in\mathbb{R}^{n}$. If $p(x)\in\Sigma[x]$, it can be expanded to
$p(x)=Z(x)^{T}QZ(x),$ (1)
@@
$\dot{x}=F_{c}(x,u)=f(x)+G(x)u,$ (2)
where $x\in\mathbb{R}^{n}$ is the state, $f(x)\in\left(R[x]\right)^{n}$ is a polynomial vector field, and $G(x)\in\left(R[x]\right)^{n\times m}$ is a polynomial matrix. System 2 with a polynomial state feedback control policy $u_{cl}(x)\in\left(R[x]\right)^{m}$ results in a closed-loop system of the form
@@
The interior of a set $\mathcal{X}\subseteq\mathbb{R}^{n}$ is denoted by $\text{int}(\mathcal{X})$. A subset $\mathcal{X}\subseteq\mathbb{R}^{n}$ is called forward invariant (cf. [15, Theorem 4.4]) with respect to system 3 if for every $x(0)\in\mathcal{X}$, $x(t)\in\mathcal{X}$ for all $t\geq 0$. A system 3 is called safe (cf. [16]) w.r.t. an allowable set of states $\mathcal{X}_{a}\subseteq\mathbb{R}^{n}$ and the safe set $\mathcal{X}_{s}\subseteq\mathbb{R}^{n}$, if $\mathcal{X}_{s}$ is forward invariant and $\mathcal{X}_{s}\subseteq\mathcal{X}_{a}$. We say that a controller $u_{cl}(x)$ guarantees safety if the corresponding closed-loop system is safe.
@@ II-B Putinar’s Positivstellensatz
The following theorem states Putinar’s P-satz (cf. [17, Theorem 3.20]).
@@
In words, if the set in 5 is empty, Putinar’s P-satz ensures that the polynomials $\gamma_{0},\gamma_{2},...,\gamma_{n}$ defined in 4 exist, albeit without specifying their degree. To find these polynomials computationally requires to iteratively increase the degree of the polynomials until a solution can be found for 6. An increase in the degree of the polynomials will, however, deteriorate computation time. For many practical examples of interest, however, low-degree polynomials suffice.
Condition 6 can also be expressed more explicitly through the condition that there exists $\gamma_{1},\gamma_{2},...,\gamma_{n}\in\Sigma[x]$ such that
@@
Assuming the empty-set condition in 5 holds, a solution to 7 involving $\gamma_{1}(x)=1$ is guaranteed by Putinar’s P-satz to exist when the degrees of the polynomials are possibly unbounded. In this case, the introduction of the SOS polynomial $\gamma_{1}(x)$ needlessly increases the set of feasible polynomials $f_{1}(x)$. However, when the degrees of the polynomials are truncated, this assertion may no longer hold true, justifying the utilization of $\gamma_{1}(x)$. For the rest of the paper, we refer to both inclusions, 6 and 7, as an SOS constraint.
@@ III-A Specifications
The concept of a safety filter (see Figure 2) is based on the idea of enhancing the behavior of a previously designed legacy controller $u_{n}(x)$ with safety guarantees. The safety filter adjusts the control action of the legacy controller, ideally in a minimal way, to render the system safe with respect to an allowable set $\mathcal{X}_{a}$, encoding the state constraints, and a safe set $\mathcal{X}_{s}$. The safe set denotes a set of states that is forward invariant for some state feedback controller. Therefore, we require that the safe set is strictly contained within the allowable set, i.e.,
@@
$\mathcal{X}_{n}\subsetneq\mathcal{X}_{s}.$ (10)
For a legacy controller that stabilizes the closed-loop system to an attractor within $\mathcal{X}_{a}$, a forward invariant set does always exist.
Moreover, the advanced safety filter should preserve such a stability guarantee across the entire safe set. This is already accomplished within the nominal region $\mathcal{X}_{n}$ due to its forward invariance property. Hence, we only need to additionally ensure a finite-time convergence to $\mathcal{X}_{n}$ for all trajectories originating within the transitional region, defined as:
@@ III-B CBF and CLF conditions formalizing the specifications
Figure 3: The advanced safety filter is characterized by the following specifications: (a) the safe set $\mathcal{X}_{s}$ is forward invariant as encoded in 15, (b) it ensures finite-time convergence towards the nominal region $\mathcal{X}_{n}$ as encoded in 17, and (c) the nominal region is forward invariant under the legacy controller $u_{n}(x)$ as encoded in 19 assuming $d(x)=0$.
We employ control Lyapunov and barrier functions to encode the specification from the previous subsection by introducing energy-like scalar functions. In the process, we define the safe set as the intersection of zero-sublevel sets of multiple continuously differentiable functions to closely fit the algebraic structure of the allowable set
$\mathcal{X}_{a}:=\Bigl{\{}x\mid w_{i}(x)\leq 0\text{ }\forall i\in\mathcal{I}_{t}\Bigr{\}}=\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{a,i},$ (12)
where $w_{i}\in R[x]$, $i\in\mathcal{I}_{t}$, are scalar polynomials, and their zero-sublevel sets $\mathcal{X}_{a,i}:=\{x\mid w_{i}(x)\leq 0\}$ are assumed to be compact. This compactness condition is important for applying Putinar’s P-satz in the following section of the paper, and it is not restrictive for the applications we have in mind.
The forward invariance property of the safe set can be asserted by a scalar function $B(x)$ defined as the point-wise maximum of a set of continuously differentiable functions $B_{i}:\mathbb{R}^{n}\to\mathbb{R}$ indexed by $\mathcal{I}_{t}$:
$B(x):=\underset{i\in\mathcal{I}_{t}}{\text{max}}\,B_{i}(x),$ (13)
@@
$\mathcal{X}_{s}:=\left\{x\mid B(x)\leq 0\right\}=\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{s,i},$
where $\mathcal{X}_{s,i}:=\left\{x\mid B_{i}(x)\leq 0\right\}$. Taking the intersection of multiple functions $B_{i}(x)$ allows for greater flexibility to increase the volume of the safe set compared to using the zero-sublevel set of a single CBF function. This is because the zero-sublevel sets of each $B_{i}(x)$ can extend beyond the allowable set, provided their intersection remains within the allowable set. In contrast, the zero-sublevel set of a single CBF must be fully contained within the allowable set. Hence, we encode the containment condition in 8 as:
@@
for every $i\in\mathcal{I}_{t}$. This can be seen from $\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{s,i}\subseteq\bigcap_{i\in\mathcal{I}_{t}}\mathcal{X}_{a,i}$ implying $\mathcal{X}_{s}\subseteq\mathcal{X}_{a}$. Here, we assume that the CBF $B(x)$ is defined using the same number of polynomials $B_{i}(x)$ as polynomials $w_{i}(x)$ that define the allowable set. To reduce the number of decision variables in $B(x)$, we can equate certain polynomials $B_{i}(x)$ in the resulting SOS problem. This reduction results in a safe set comprising fewer distinct subsets $\mathcal{X}_{s,i}$ compared to the subsets $\mathcal{X}_{a,i}$ that define the allowable set.
@@
We assume the safe set $\mathcal{X}_{s}$ to be regular in the sense that $\forall x\in\mathcal{X}_{s},\exists z\in\mathbb{R}^{n},\forall i\in\mathcal{I}_{t}$ [18, Definition 4.9]:
$B_{i}(x)+\nabla B_{i}(x)^{T}z<0$
The regularity assumption in Assumption 2 is a mild one, and it generically holds true, if $B_{i}(x)$ is chosen as a polynomial. It implies that the interior of the safe set is given by $\text{int}(\mathcal{X}_{s})=\left\{x\mid B_{i}(x)<0\quad i\in\mathcal{I}_{t}\right\}$. Examples of polynomials that violate the regularity assumption can be found in [18, Chapter 4.2].
To ensure forward invariance of $\mathcal{X}_{s}$, as shown in Figure 3(a), $B(x)$ needs to satisfy Nagumo’s theorem, that is, $\forall x\in\partial\mathcal{X}_{s},\exists u\in\mathbb{R}^{m}$ (cf. [18, Nagumo’s Theorem 4.7]):
@@
$\mathcal{T}_{\mathcal{X}_{s}}(x):=\{z\mid\nabla B_{i}(x)^{T}z\leq 0\quad i\in\text{Act}(x)\},$ (16)
and $\text{Act}(x):=\left\{i\in\mathcal{I}_{t}\mid B_{i}(x)=0\right\}$ denotes the set of active constraints. The function $B(x)$ in 13 satisfying 15 and 16 will henceforth be referred to as a Control Barrier Function (CBF).
Compared to the standard safety filter, our approach also require finite-time convergence towards $\mathcal{X}_{n}$, as outlined in the specifications. This is asserted by a Control Lyapunov Function (CLF) defined by a continuously differentiable function $V:\mathbb{R}^{n}\to\mathbb{R}$. This function needs to satisfy a Control Lyapunov-like condition in the transitional region $\mathcal{X}_{t}$, as shown in Figure 3(b). Namely, $\forall x\in\mathcal{X}_{t},\exists u\in\mathbb{R}^{m}$:
$\nabla V(x)^{T}\left(f(x)+G(x)u\right)+d(x)\leq 0$ (17)
@@
where $\lambda:\mathbb{R}^{n}\to\mathbb{R}_{>0}$ is a Lipschitz continuous function, and $V_{0}\leq\text{min}_{x}\left(V(x)\right)$ is a constant value, ensuring $d(x)$ is strictly positive (and thus $V(x)$ strictly decaying) in the transitional region $\mathcal{X}_{t}$. By utilizing a strictly positive function $d(x)$, the CLF condition can be expressed as a non-strict inequality condition similar to the CBF condition – a crucial distinction for the subsequent arguments presented in the next section. We define the nominal region as the zero sublevel set of $V(x)$:

### v3

@@ Abstract
This paper presents a novel safety filter framework that ensures both safety and the preservation of the legacy control action within a nominal region. This modular design allows the safety filter to be integrated into the control hierarchy without compromising the performance of the existing legacy controller during nominal operation. For a control-affine system, this is accomplished by formulating multiple Control Barrier Functions (CBFs) and Control Lyapunov-like Functions (CLFs) conditions, alongside a forward invariance condition for the legacy controller, as sum-of-squares constraints. Additionally, the state-dependent inequality constraints of the resulting Quadratic Program (QP) – encoding the CBF and CLF conditions – are designed to remain inactive within the nominal region, ensuring preservation of the legacy control action and performance. Our safety filter design is also the first to include quadratic input constraints, and does not need an explicit specification of the attractor, as it is implicitly defined by the legacy controller. To avoid the chattering effect and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the QP are selected to be regular. Finally, we demonstrate the method in a detailed case study involving the control of a three-phase ac/dc power converter.
@@ I Introduction
In several engineering applications, it is of much interest to endow a well-performing controller with safety guarantees by minimally changing its control action. This concept is commonly referred to as a safety filter, and a comprehensive overview of the subject can be found in [21]. An illustrative example is a finely-tuned control design based on linearization, which must be extended to adhere to hard state constraints in the nonlinear closed-loop system. We refer to such a controller as the legacy controller. A popular type of safety filters for control-affine systems relies on a Control Barrier Function (CBF). By continuously solving the associate Quadratic Program (QP), the control action of the legacy controller is smoothly adjusted as the system approaches the boundary of the safe set $\mathcal{X}_{s}$. The main challenges in deploying such a safety filter lies in finding viable CBF candidates and adhering to input constraints.
Even though safety filters based on a QP are well-established, they are frequently not adopted in current industrial setups due to their risk of altering the legacy control behavior during nominal operations. In practice, the legacy control parameters are often meticulously tuned and thoroughly evaluated, which require significant time and resources. As a result, preserving the legacy control action within the nominal operating regime is not merely a preference but a strict engineering requirement. In such contexts, the role of a safety filter is to act as a last-resort mechanism that intervenes only when the system state approaches critical boundaries.
@@
This paper presents a safety filter concept that guarantees the preservation of the legacy control action within a nominal region $\mathcal{X}_{n}$, which contains the desired closed-loop equilibrium point (or another attractor), as depicted in Figure 1. Unlike in a reach-avoid problem [11], the nominal region $\mathcal{X}_{n}$ does not represent a predetermined set of target states. Instead, it is selected during the design process (often at an earlier stage) to encompass the nominal operating regime. The specifications are inherently more stringent than those of a standard safety filter, but may therefore be deployed in an industrial setup without disrupting the original control architecture.
Our safety filter is realized for control-affine systems based on solving a QP at every state, where the linear inequality constraints are derived from multiple CBFs, Control Lyapunov-like Functions (CLFs), and state-dependent slack variables. These slack variables provide local flexibility by relaxing the constraint in regions where the CBF and CLF condition are not required to hold. A Quadratic Constraint Quadratic Program (QCQP) instead of a simple QP can be employed to encode quadratic input constraints, which are common, e.g., in power electronic applications. Despite the restriction to control-affine and polynomial systems, our approach remains relevant for a wide range of systems encountered in practice, especially in domains like power systems and robotics, where control-affine polynomial models are commonly used. The design process consists of two stages: first, we identify compatible CBF and CLF candidates using a Sum-of-Squares (SOS) optimization approach; second, we determine the state-dependent slack variables via a separate SOS program, ensuring that the legacy control action is preserved within the nominal region. Our approach focuses on computing formal guarantees via SOS optimization, enabling an automated design procedure and eliminating the need for application-specific parameter tuning based on extensive simulations.
@@
The proposed safety filter concept resembles the integration of a CLF with a CBF to attain both safety and stabilization. Such an integration, although just with a single function, has been initially proposed in [17], and it has since become a well-established concept, see the survey in [1]. Feasibility of such an optimization-based controller can be accomplished by introducing a slack variable that deactivates CLFs and CBFs based on their respective priorities, as shown in [2]. Alternatively, it is possible to compute compatible CLFs and CBFs using SOS tools, ensuring that for every state, there exists an input $u$ satisfying all CBF and CLF conditions. In order to encode compatibility, there is one approach that introduces an SOS constraint for each CLF/CBF pair in the absence of inputs constraints [6]. The second approach establishes compatibility by introducing a polynomial or rational state-feedback controller, as outlined in [23, 18]. Such a controller can also be used to enforce linear input constraints, as demonstrated in [22]. Another noteworthy approach for imposing input constraints using a contrapositive statement, albeit without ensuring compatibility between CBFs and CLFs, is introduced in [7]. These approaches introduce auxiliary polynomials that render the resulting SOS problem bilinear in its decision variables. In [19], an alternating algorithm is proposed to solve this bilinear SOS problem. One challenge in seeking CLF and CBF candidates through an alternating algorithm is to identify an initial feasible point. In their work in [25], the authors constrained their search for an initial CBF to scalar functions of a specific form. This restriction allows to reformulate the problem with linearity in its decision variables.
After our previous work [18] on synthesizing compatible CLF and multiple CBFs using SOS optimization, we identified a recurring issue: industrial standards often requires compliance with pre-certified control behavior, particularly during nominal operations. A prime example is the integration of a safety filter. To address this, we proposed a modular and minimal invasive safety filter approach that preserves the legacy control behavior within the nominal region, while still ensuring safety near constraint boundaries. To address those issues we extended the framework by including the following contributions:
  • First, we synthesize compatible CLFs and CBFs based on an SOS problem, augmented by an additional SOS constraint that ensures forward-invariance of the nominal region under the legacy control. This constraint is essential, as it enables the construction of the proposed safety filter in a subsequent step.
@@
  • Second, to increase flexibility in shaping the safe set, we employ multiple CBFs. Rather than imposing each CBF condition separately, we relax these conditions when the corresponding CBF is not active, similar to the idea in [8]. Moreover, the CLF condition can also be relaxed by leveraging the forward invariance property of the CBFs, similar to [20] for autonomous systems.
  • Third, given feasible CLF and CBFs candidates, we formulate a second SOS problem to compute the state-dependent slack variables that ensure safety while preserving the legacy behavior within the nominal region.
@@
  • Fourth, we introduce a novel approach for encoding linear and quadratic input constraints on the rational controller. This ensures feasibility of the resulting QP-based controller, even in case of quadratic input constraints – an issue that, to the authors’ knowledge, is unsolved.
@@
To avoid chattering and guarantee the uniqueness and Lipschitz continuity of solutions, the state-dependent inequality constraints of the QP are selected to be regular (see Assumption 2). Finally, we tested our framework on a detailed three-phase ac/dc converter system connected to an infinite bus, where we successfully synthesized a safety filter that ensures satisfaction of the state constraints given by the dc-link voltage and line current, while adhering to input constraints.
@@ II-A Notation
The shorthand $\mathcal{I}_{n}:=\{1,2,...,n\}$ designates a range of natural numbers from $1$ to $n$. $R[x]$ refers to the set of scalar polynomials in variables $x\in\mathbb{R}^{n}$, and $\Sigma[x]$ refers to the set of scalar SOS polynomials in variables $x\in\mathbb{R}^{n}$. If $p(x)\in\Sigma[x]$, it can be expanded to
$p(x)=Z(x)^{\top}QZ(x),$ (1)
@@
$\dot{x}=f(x)+G(x)u,$ (3)
where $x\in\mathbb{R}^{n}$ is the state, $f(x)\in\left(R[x]\right)^{n}$ is a polynomial vector field, and $G(x)\in\left(R[x]\right)^{n\times m}$ is a polynomial matrix. System 3 with a polynomial state feedback control policy $u_{cl}(x)\in\left(R[x]\right)^{m}$ results in a closed-loop system of the form
@@
The interior of a set $\mathcal{X}\subseteq\mathbb{R}^{n}$ is denoted by $\text{int}(\mathcal{X})$. A subset $\mathcal{X}\subseteq\mathbb{R}^{n}$ is called forward invariant (cf. [10, Theorem 4.4]) with respect to system 4 if for every $x(0)\in\mathcal{X}$, $x(t)\in\mathcal{X}$ for all $t\geq 0$. A system 4 is called safe (cf. [24]) w.r.t. an allowable set of states $\mathcal{X}_{a}\subseteq\mathbb{R}^{n}$ and the safe set $\mathcal{X}_{s}\subseteq\mathbb{R}^{n}$, if $\mathcal{X}_{s}$ is forward invariant and $\mathcal{X}_{s}\subseteq\mathcal{X}_{a}$. In the following, the allowable set $\mathcal{X}_{a}$ is assumed compact and regular. We say that a controller $u_{cl}(x)$ guarantees safety if the corresponding closed-loop system is safe.
@@ II-B Putinar’s Positivstellensatz
The following theorem states Putinar’s P-satz (cf. [13, Theorem 3.20]).
@@
In words, if the set in 6 is empty, Putinar’s P-satz ensures that the polynomials $\gamma_{0},\gamma_{2},...,\gamma_{n}$ defined in 5 exist, albeit without specifying their degree. To find these polynomials computationally requires to iteratively increase the degree of the polynomials until a solution can be found for 7. An increase in the degree of the polynomials will, however, deteriorate computation time. For many practical examples of interest, however, low-degree polynomials suffice.
Condition 7 can also be expressed more explicitly through the condition that there exists $\gamma_{1},\gamma_{2},...,\gamma_{n}\in\Sigma[x]$ such that
@@
Assuming the empty-set condition in 6 holds, a solution to 8 involving $\gamma_{1}(x)=1$ is guaranteed by Putinar’s P-satz to exist when the degrees of the polynomials are possibly unbounded. In this case, the introduction of the SOS polynomial $\gamma_{1}(x)$ needlessly increases the set of feasible polynomials $f_{1}(x)$. However, when the degrees of the polynomials are truncated, this assertion may no longer hold true, justifying the utilization of $\gamma_{1}(x)$. For the rest of the paper, we refer to both inclusions, 7 and 8, as an SOS constraint.
@@ III-A Specifications
The main objective of a safety filter (see Figure 2) is to enhance the behavior of a previously designed legacy controller $u_{n}(x)$ with safety guarantees. The safety filter adjusts the control action of the legacy controller, ideally in a minimal way, to ensure the system remains within a safe set $\mathcal{X}_{s}$, which is contained within an allowable set $\mathcal{X}_{a}$ that encodes the state constraints:
@@
$\mathcal{X}_{n}\subseteq\mathcal{X}_{s}.$ (11)
For a legacy controller that stabilizes the closed-loop system to an attractor within $\mathcal{X}_{a}$, such a forward invariant set does always exist.
Moreover, the advanced safety filter is designed to preserve the attractor of the legacy controller – assuming it lies within $\mathcal{X}_{a}$. This property is automatically satisfied within $\mathcal{X}_{n}$ due to the preservation property 10 and forward invariance of $\mathcal{X}_{n}$. Thus, it remains to ensure finite-time convergence of all trajectories originating outside $\mathcal{X}_{n}$ but within the safe set $\mathcal{X}_{s}$. We define the corresponding transitional region as:
@@ III-B CBF and CLF conditions formalizing the specifications
Figure 3: The advanced safety filter is characterized by the following specifications: (a) the safe set $\mathcal{X}_{s}$ is forward invariant as encoded in 16, (b) it ensures finite-time convergence towards the nominal region $\mathcal{X}_{n}$ as encoded in 18, and (c) the nominal region is forward invariant under the legacy controller $u_{n}(x)$ as encoded in 20 assuming $d(x)=0$.
We employ control Lyapunov and barrier functions to encode the specification from the previous subsection by introducing energy-like scalar functions. In the process, we define the safe set as the intersection of zero-sublevel sets of multiple continuously differentiable functions
$\mathcal{X}_{a}:=\Bigl\{x\mid w_{i}(x)\leq 0\quad\forall i\in\mathcal{I}_{n_{\!B}}\Bigr\}=\bigcap_{i\in\mathcal{I}_{n_{\!B}}}\mathcal{X}_{a,i},$ (13)
where $w_{i}\in R[x]$ are scalar polynomials, and $\mathcal{X}_{a,i}:=\{x\mid w_{i}(x)\leq 0\}$ denotes the corresponding zero-sublevel set. We assume that each set $\mathcal{X}_{a,i}$ is compact, which ensures that their intersection $\mathcal{X}_{a}$ is also compact. This compactness condition is important for the derivation of the SOS constraints, and it is well aligned with our intended application scenarios, where state constraints are typically bounded.
The forward invariance property of the safe set can be asserted by a scalar function $B(x)$ defined as the point-wise maximum of a set of continuously differentiable functions $B_{i}:\mathbb{R}^{n}\to\mathbb{R}$ indexed by $\mathcal{I}_{n_{\!B}}$:
$B(x):=\underset{i\in\mathcal{I}_{n_{\!B}}}{\text{max}}\,B_{i}(x),$ (14)
@@
$\mathcal{X}_{s}:=\left\{x\mid B(x)\leq 0\right\}=\bigcap_{i\in\mathcal{I}_{n_{\!B}}}\mathcal{X}_{s,i},$
where $\mathcal{X}_{s,i}:=\left\{x\mid B_{i}(x)\leq 0\right\}$. This multi-function construction provides greater flexibility in shaping the safe set compared to relying on a single barrier function. In particular, each function $B_{i}(x)$ is allowed to define a sublevel set that may partially extend beyond the allowable set, as long as their intersection $\mathcal{X}_{s}$ remains within the allowable set $\mathcal{X}_{a}$. By contrast, a single barrier function must define a sublevel set that is entirely contained within the allowable set, which leads to more conservative safe sets. Hence, to enforce the containment condition in 9, we impose the subset relation
@@
for every $i\in\mathcal{I}_{n_{\!B}}$. This condition ensures that the overall safe set $\mathcal{X}_{s}=\bigcap_{i\in\mathcal{I}_{n_{\!B}}}\mathcal{X}_{s,i}$ is contained within the allowable set $\mathcal{X}_{a}=\bigcap_{i\in\mathcal{I}_{n_{\!B}}}\mathcal{X}_{a,i}$. We assume that the barrier function $B(x)$ is constructed from the same number of components $B_{i}(x)$ as the number of polynomials $w_{i}(x)$ that define the allowable set. This condition is not restrictive in practice, since one can define multiple $B_{i}(x)$ functions to share the same expression, effectively reducing the number of distinct subsets $\mathcal{X}_{s,i}$ compared to the subsets $\mathcal{X}_{a,i}$. Conversely, it is also possible to define additional distinct $B_{i}(x)$ functions even if some $w_{i}(x)$ are repeated.
@@
We assume the safe set $\mathcal{X}_{s}$ to be regular in the sense that $\forall x\in\mathcal{X}_{s},\exists z\in\mathbb{R}^{n},\forall i\in\mathcal{I}_{n_{\!B}}$ [5, Def. 4.9]:
$B_{i}(x)+\nabla B_{i}(x)^{\top}z<0$
The regularity assumption in Assumption 2 is a mild one, and it generically holds true, if $B_{i}(x)$ is chosen as a polynomial. It implies that the interior of the safe set is given by $\text{int}(\mathcal{X}_{s})=\left\{x\mid B_{i}(x)<0\,\,\forall i\in\mathcal{I}_{n_{\!B}}\right\}$. Examples of polynomials that violate the regularity assumption can be found in [5, Chapter 4.2].
To ensure forward invariance of $\mathcal{X}_{s}$, as shown in Figure 3(a), $B(x)$ needs to satisfy Nagumo’s theorem, that is, $\forall x\in\partial\mathcal{X}_{s},\exists u\in\mathbb{R}^{m}$ (cf. [5, Nagumo’s Theorem 4.7]):
@@
$\mathcal{T}_{\mathcal{X}_{s}}(x):=\{z\mid\nabla B_{i}(x)^{\top}z\leq 0\quad i\in\text{Act}(x)\},$ (17)
and $\text{Act}(x):=\left\{i\in\mathcal{I}_{n_{\!B}}\mid B_{i}(x)=0\right\}$ denotes the set of active constraints. The function $B(x)$ in 14 satisfying 16 and 17 will be referred to as a Control Barrier Function (CBF).
Compared to the standard safety filter, our approach additionally require finite-time convergence to the nominal region $\mathcal{X}_{n}$. This is enforced via a Control Lyapunov Function (CLF), defined by a continuously differentiable function $V:\mathbb{R}^{n}\to\mathbb{R}$. In the transitional region $\mathcal{X}_{t}$, this function needs to satisfy a Control Lyapunov-like condition, as illustrated in Figure 3(b). Namely, $\forall x\in\mathcal{X}_{t},\exists u\in\mathbb{R}^{m}$:
$\nabla V(x)^{\top}\left(f(x)+G(x)u\right)+d(x)\leq 0$ (18)
@@
where $\lambda:\mathbb{R}^{n}\to\mathbb{R}_{>0}$ is a Lipschitz continuous function, and $V_{0}\leq\text{min}_{x}\left(V(x)\right)$ is a constant value, ensuring $d(x)$ is strictly positive (and thus $V(x)$ strictly decaying) in the transitional region $\mathcal{X}_{t}$. By utilizing a strictly positive function $d(x)$, the CLF condition can be expressed as a non-strict inequality condition similar to the CBF condition – an important distinction for deriving the SOS constraints. We define the nominal region as the zero sublevel set of $V(x)$: